# DesCLIP: Robust Continual Adaptation via General Attribute Descriptions for Pretrained Vision-Language Models

### v1

@@ III-A Preliminaries
1) Continual Learning Formulation: A sequence of task datasets is denoted as $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{T}\}$. During training on task $\mathcal{D}_{t}$ ($t\in\{1,2,\ldots,T\}$), access to data from previous tasks $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{t-1}\}$ is either highly restricted or entirely unavailable. In class-incremental learning (CIL), datasets for different tasks are introduced sequentially. Each task $t$ is associated with a unique set of classes $\mathcal{C}_{t}=\{C_{t,1},C_{t,2},\ldots,C_{t,|\mathcal{C}_{t}|}\}$, where $|\mathcal{C}_{t}|$ denotes the number of classes in task $t$. The classes associated with different tasks are disjoint:
$\mathcal{C}_{t}\cap\mathcal{C}_{t^{\prime}}=\emptyset,\quad\forall t\neq t^{\prime},\,t,t^{\prime}\in\{1,2,\ldots,T\}.$ (1)
@@
Here, $\mathbf{RP}_{y}$ is derived from hand-crafted prompts, typically following a template such as “A photo of a [CLS],” where [CLS] represents the specific class name. The probability of classifying a test image $x$ as class $y_{i}$ is computed using the softmax function:
@@
$p\left(y_{i}\mid x\right)=\frac{\exp\left(\left\langle\frac{\mathbf{z}}{\|\mathbf{z}\|},\frac{\mathbf{w}_{y_{i}}}{\|\mathbf{w}_{y_{i}}\|}\right\rangle/\tau\right)}{\sum_{\mathbf{w}_{j,k}\in\mathbf{W}_{t}}\exp\left(\left\langle\frac{\mathbf{z}}{\|\mathbf{z}\|},\frac{\mathbf{w}_{j,k}}{\|\mathbf{w}_{j,k}\|}\right\rangle/\tau\right)}.$ (5)
@@ III-C General Attribute Description Generation
CLIP establish robust visual-textual associations during the pre-training phase through instance-level contrastive learning. However, most existing research overlooks this foundational capability, conventionally relying on fixed, hand-crafted templates combined with class names as prompts to derive “prior” classification weights via the textual encoder. Although Wang et al. [16] introduced an “attribute bank” to enable attribute sharing across different task instances, this approach lacks intrinsic relevance to specific classes. For instance, attributes such as “white” and “grass” fail to provide meaningful distinctions between classes like “cat” and “dog”.
@@
  • Q: Describe what does a/an [CLS] look like?
  • Q: Describe a/an [CLS]’s attribute features.
  • Q: Describe a/an [CLS]’s outlook features.
@@
  • Q: Describe what are some attribute characteristics of [CLS] compared with other [P-CLS], visually.
  • Q: Describe what kind of [P-CLS] a/an [CLS] is, visually.
Here, [P-CLS] refers to the parent class of [CLS]. Fine-grained DRPs are also employed for tasks with a known general scope, such as identifying objects within the “birds” supercategory:
  • Describe a/an [CLS]’s attributes from its beak, eyes, body, belly, tail, wings, breast, etc.
The DRP-guided general attribute description generation are illustrated in Fig. 4. The language assistant generates $n_{dsc}$ attribute description candidates (DCs) for the $k$-th class of incremental task $t$, denoted as:
@@ III-D Anchor-based Embedding Filter
[…]
where $j\in\{1,2,\dots,n_{dsc}\}$. To mitigate the risk of overfitting in the CLIP’s visual encoder, visual features with low relevance to either class text or GA descriptions should be filtered out. Hence, a condition $\chi(\mathbf{z}_{i})$ is defined to filter visual features as:
@@
We posit that embedding candidates in $\mathbf{EC}_{c}$ exhibiting a similarity score surpassing a threshold $\gamma$ (relative to the anchor threshold $AT_{i}^{c}$) are more consistent with the visual features of the current sample. These candidates are filtered as follows:
@@
[…]
@@ III-E General Attribute-Guided Progressive Visual-Textual Alignment
1) Instance Matching: To align the visual features with highly relevant embedded GA descriptions, we select the most closest textual representation $\mathbf{h}_{i}$ as the paired text embedding:
@@
$\mathcal{L}_{\text{IM}}=\mathbb{E}_{i\in P}\Big{[}-\log\frac{\mathbf{M}_{i,i}}{\mathbf{M}_{i,i}+\sum_{j\in P,j\neq i}\mathbf{M}_{i,j}}\Big{]},$ (14)
@@
TABLE I: Comparison of different methods on the CIFAR100 [29], ImageNet-Subset [24], and CUB-200 [30] datasets under various settings. ‘UB’ denotes the upper bound achieved through joint training.
Method CIFAR100 [29] ImageNet-Subset [24] CUB-200 [30]
$T=5$ $T=10$ $T=20$ $T=5$ $T=10$ $T=20$ $T=5$ $T=10$ $T=20$
Seq FT - - 50.6 66.9 - - - - 58.8 73.8 - - - - 40.8 64.6 - -
LwF [11] - - 56.1 74.8 - - - - 56.4 75.0 - - - - 52.1 71.5 - -
EWC [31] - - 64.4 78.4 - - - - 70.3 80.2 - - - - 64.4 78.6 - -
ContinualCLIP [58] 72.8 - 72.8 - 72.8 - 74.6 - 74.6 - 74.6 - 60.9 - 60.9 - 60.9 -
L2P [44] - - 70.2 79.6 - - - - 71.1 78.4 - - - - 69.9 77.8 - -
DualPrompt [44] - - 72.0 81.8 - - - - 71.7 79.8 - - - - 64.5 75.9 - -
SLCA [59] 80.2 87.2 80.2 87.6 80.2 87.6 82.3 86.0 80.3 84.4 80.2 85.4 75.5 78.9 73.9 80.0 71.1 79.4
UB 90.0 - 90.0 - 90.0 - 87.6 - 87.6 - 87.6 - 84.5 - 84.5 - 84.5 -
TABLE II: Comparison of various methods on CIFAR100 [29] and CUB-200 [30] ($T=10$) with CLIP ViT-B/16 [1] backbone. ‘#FR’ represents using feature replay. ‘#AS’ represents using additional structure.
Method CIFAR100 [29] CUB-200 [30]
#FR #AS Last Avg C. Last Avg C.
Seq FT 46.3 - 24.2 45.7 - 39.7
ContinualCLIP[58] 68.3 - 63.6 55.1 - 63.6
L2P [44] 64.5 72.9 41.8 66.3 75.5 43.8
SLCA [59] 71.5 79.8 59.1 50.6 58.4 62.7
AttriCLIP [16] 67.0 77.8 60.3 50.8 65.4 63.1
SPU [10] 75.8 84.3 58.7 67.8 75.6 59.8
TaskRes-CL [5] 75.6 83.2 63.6 67.1 73.8 63.6
MoE-Adapter [18] 77.8 84.9 62.2 68.3 76.2 62.9
RAPF [60] 78.5 85.1 54.0 73.2 80.4 52.6
DesCLIP (Ours) 79.1 85.7 62.0 72.0 78.6 62.5
$\mathcal{L}_{\text{TA}}^{i}=\mathbb{E}_{\mathbf{u}\in\mathbf{FE}^{i}}\Big{[}\beta\cdot\Big{\|}\frac{\mathbf{w}^{\prime}_{t,k}}{\|\mathbf{w}^{\prime}_{t,k}\|}-\frac{\mathbf{u}}{\|\mathbf{u}\|}\Big{\|}_{2}+\Big{(}1-\left\langle\frac{\mathbf{w}^{\prime}_{t,k}}{\|\mathbf{w}^{\prime}_{t,k}\|},\frac{\mathbf{u}}{\|\mathbf{u}\|}\right\rangle\Big{)}\Big{]},$ (19)
where $\beta$ is a parameter that determines whether the alignment of text embeddings places greater emphasis on their absolute distance in the text space (e.g., Euclidean distance) or on the directional consistency. Text alignment loss $\mathcal{L}_{\text{TA}}$ of the current batch is: $\mathcal{L}_{\text{TA}}=\sum_{i\in P}\mathcal{L}_{\text{TA}}^{i}.$
Since, in the context of continual learning, data from previous tasks cannot be revisited during the current task, the weights $\mathbf{w^{\prime}}_{t}=\{\mathbf{w^{\prime}}_{t,1},\mathbf{w^{\prime}}_{t,2},...,\mathbf{w^{\prime}}_{t,|C_{t}|}\}$ are calibrated solely within the scope of the current task $t$. Consequently, only the shifting weights $\mathbf{s}_{t}=\{\mathbf{s}_{t,1},\mathbf{s}_{t,2},...,\mathbf{s}_{t,|C_{t}|}\}$ associated with the current task are learnable, whereas the shifting weights $\mathbf{s}_{0:t-1}=\{\mathbf{s}_{0},\mathbf{s}_{1},...,\mathbf{s}_{t-1}\}$ from previous tasks remain fixed.
@@
$\mathcal{L}_{\text{RIC}}=\mathbb{E}_{i\in\{1,2,...,B\}}\Bigg{[}\log\frac{-\exp\Big{(}\left\langle\frac{\mathbf{z}_{i}}{\|\mathbf{z}_{i}\|},\frac{\mathbf{w^{\prime}}_{y_{i}}}{\|\mathbf{w^{\prime}}_{y_{i}}\|}\right\rangle/\tau\Big{)}}{\sum_{k=1}^{|C_{t}|}\exp\Big{(}\left\langle\frac{\mathbf{z}_{i}}{\|\mathbf{z}_{i}\|},\frac{\mathbf{w^{\prime}}_{t,k}}{\|\mathbf{w^{\prime}}_{t,k}\|}\right\rangle/\tau\Big{)}}\Bigg{]},$ (20)
@@ III-F Inference Stage
The GA descriptions and embeddings we introduce do not participate in the inference stage of the CLIP after each training phase, which effectively avoids additional storage overhead and eliminates any increase in inference time. Consequently, the model obtained through our method maintains identical parameter size and inference time during the inference stage as the original zero-shot CLIP. In inference stage after task $t$, we leverage the calibrated text embeddings of all seen classes: $\mathbf{W^{\prime}}_{t}=\bigcup_{j=1}^{t}\bigcup_{k=1}^{|C_{j}|}\mathbf{w^{\prime}}_{j,k}.$ Hence, the probability of predicting the testing image $x$ as the class $y_{i}$ can be expressed as:
$p\left(y_{i}\mid x\right)=\frac{\exp\left(\left\langle\frac{\mathbf{z}}{\|\mathbf{z}\|},\frac{\mathbf{w^{\prime}}_{y_{i}}}{\|\mathbf{w^{\prime}}_{y_{i}}\|}\right\rangle/\tau\right)}{\sum_{\mathbf{w}^{\prime}_{j,k}\in\mathbf{W^{\prime}}_{t}}\exp\left(\left\langle\frac{\mathbf{z}}{\|\mathbf{z}\|},\frac{\mathbf{w}^{\prime}_{j,k}}{\|\mathbf{w}^{\prime}_{j,k}\|}\right\rangle/\tau\right)}.$ (22)
@@ IV-A Setup
Figure 5: Average accuracy evaluated after each task across different incremental task settings (5, 10, and 20 tasks).
Figure 6: Degradation or improvement evaluation on ImageNet-Subset [24] with 10 incremental tasks. Upper: accuracy differences relative to zero-shot VLM. Bottom: accuracy differences relative to VLM previously learned on the first-half classes.
Figure 7: Performance comparison in few-shot case.
Figure 8: Ablation study of generated GA description amount.
1) Datasets: The evaluation experiments for continual learning are conducted on CIFAR100 [29], ImageNet-Subset [24] and CUB-200 [30]. ImageNet-full [24] is utilized as a control set to evaluate the retention of the pretrained generalization knowledge in CLIP.
  • CIFAR100 [29] introduced a dataset comprising 60,000 images, each with a resolution of $32\times 32$, distributed across 100 distinct classes. Each class includes 500 training samples and 100 testing samples. CIFAR100 has become a widely recognized benchmark for evaluating continual learning methods. Its low-resolution images pose a significant challenge for VLMs that are pre-trained on high-resolution datasets, highlighting the difficulty of adapting to such datasets in continual learning scenarios.
  • ImageNet-Subset [24]. Following the class configuration proposed in [32], we select the challenging ImageNet-Subset benchmark for method evaluation. This subset consists of fine-grained animal images from 100 categories that are unfamiliar to VLMs and prone to misclassification due to their subtle differences.
  • Caltech-UCSD Birds-200 (CUB-200) [30] comprises 11,788 bird images distributed across 200 categories. The subtle variations among bird images pose significant challenges for VLMs to achieve accurate identification. Moreover, the relatively small size of the dataset further intensifies the difficulty, particularly in few-shot scenarios.
  • ImageNet-full [24] is a large-scale image dataset containing over 1.2 million images across 1,000 classes, widely used for pretraining and evaluating visual representation learning models.
@@
2) Metrics: To evaluate the continual learning performance of classification models, we employ two primary metrics: ‘Last’ and ‘Avg’. ‘Last’ denotes the average accuracy across all classes after the model has completed training on the final task. ‘Avg’ represents the mean incremental accuracy calculated over all tasks the model has learned thus far. In addition, The control set accuracy ‘C.’ [10] is employed to evaluate the retention of CLIP’s zero-shot generalization knowledge after continual learning, assessed on ImageNet-full [24] (1000 classes in total).
@@
3) Competitors: We compare our method against baseline and state-of-the-art methods for continual learning VLMs. These include ContinualCLIP [58] under zero-shot conditions; non-continual adapting methods such as CoOp [3] and CoOpOp [4]; and conventional continual learning method such as LwF [11] and EWC [31]. Additionally, we evaluate VLM-specific continual learning methods, including AttriCLIP [16], TaskRes-CL [5], SPU [10], MoE-Adapter [18], and RAPF [60]. For a broader comparison, we also incorporate techniques tailored to address continual learning in visual-only pre-trained models, such as L2P [44], DualPrompt [45], and SLCA [59].
@@
4) Implementation Details: All experiments are conducted on two NVIDIA GeForce RTX 3090 GPUs. Following [16], we adopt the pre-trained CLIP ViT-L/14 [1] as the backbone. To ensure a fair comparison, no data replay strategies are used during the continual learning process of the VLM.
@@
The model is trained for 10 epochs on the datasets for each incremental task. Stochastic Gradient Descent (SGD) [61] is employed as the optimizer, using a cosine learning rate decay schedule with a batch size of 32. The learning rate for the MLPs’ weights in the CLIP’s visual encoder is set to $1\times 10^{-5}$, while the learning rate for the text shifting weights is set to 0.1. The coefficient $\alpha$ for shifting weight addition is set to 0.1. Loss balancing factors are set as $\lambda_{\text{IM}}=2.0$, $\lambda_{\text{TA}}=0.5$, and $\lambda_{\text{RIC}}=1.0$ by default. For the fine-grained CUB-200 dataset [30], $\lambda_{\text{IM}}$ is adjusted to 15.0. To filter out visual features that fail to intuitively reflect attribute or class information, $\delta_{d},\gamma$ are estimated based on downstream tasks and set to 0.20, 0.015; 0.25, 0.03; and 0.30, 0.015 for CIFAR100 [29], ImageNet-Subset [24], and CUB-200 [30], respectively. Additionally, we generate $n_{dsc}=30$ attribute descriptions for each class using the language assistant GPT-4 [50].

### v2

@@ III-A Preliminaries
1) Continual Learning Formulation: A sequence of task datasets is denoted as $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{T}\}$. During training on task $\mathcal{D}_{t}$ ($t\in\{1,2,\ldots,T\}$), access to data from previous tasks $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{t-1}\}$ is either highly restricted or entirely unavailable. In class-incremental learning (CIL), datasets for different tasks are introduced sequentially. Each task $t$ is associated with a unique set of classes ${C}_{t}=\{{C}_{t,1},{C}_{t,2},\ldots,{C}_{t,|{C}_{t}|}\}$, where $|{C}_{t}|$ denotes the number of classes in task $t$. The classes associated with different tasks are disjoint:
${C}_{t}\cap{C}_{t^{\prime}}=\emptyset,\quad\forall t\neq t^{\prime},\,t,t^{\prime}\in\{1,2,\ldots,T\}.$ (1)
@@
Here, $\mathbf{RP}_{y}$ is derived from hand-crafted prompts, typically following a template such as “A photo of a <class name>.” The probability of classifying a test image $x$ as class $y_{i}$ is computed using the softmax function:
@@
$p\left(y_{i}\mid x\right)=\frac{\exp\left(\left\langle\frac{\mathbf{z}}{\|\mathbf{z}\|},\frac{\mathbf{w}_{y_{i}}}{\|\mathbf{w}_{y_{i}}\|}\right\rangle/\tau\right)}{\sum_{\mathbf{w}_{j,k}\sim\mathbf{W}_{t}}\exp\left(\left\langle\frac{\mathbf{z}}{\|\mathbf{z}\|},\frac{\mathbf{w}_{j,k}}{\|\mathbf{w}_{j,k}\|}\right\rangle/\tau\right)}.$ (5)
@@ III-C General Attribute Description Generation
CLIP establishes robust visual-textual associations during the pre-training phase through instance-level contrastive learning. However, most existing research overlooks this foundational capability, conventionally relying on fixed, hand-crafted templates combined with class names as prompts to derive “prior” classification weights via the textual encoder. Although Wang et al. [16] introduced an “attribute bank” to enable attribute sharing across different task instances, this approach lacks intrinsic relevance to specific classes. For instance, attributes such as “white” and “grass” fail to provide meaningful distinctions between classes like “cat” and “dog”.
@@
  • Q: Describe what does a/an <class name> look like?
  • Q: Describe a/an <class name>’s attribute features.
  • Q: Describe a/an <class name>’s outlook features.
@@
  • Q: Describe what are some attribute characteristics of <class name> compared with other <P-class name>, visually.
  • Q: Describe what kind of <P-class name> a/an <class name> is, visually.
Here, <P-class name> refers to the parent class of <classname><class~name>. Fine-grained DRPs are also employed for tasks with a known general scope, such as identifying objects within the “birds” supercategory:
  • Describe a/an <class name>’s attributes from its beak, eyes, body, belly, tail, wings, breast, etc.
The DRP-guided general attribute description generation is illustrated in Fig. 4. The language assistant generates $n_{dsc}$ attribute description candidates (DCs) for the $k$-th class of incremental task $t$, denoted as:
@@ III-D Anchor-based Embedding Filter
$\textbf{EC\_S}_{c}^{i,j}=\left\langle\frac{\mathbf{z}_{i}}{\|\mathbf{z}_{i}\|},\frac{\mathbf{EC}_{c,j}}{\|\mathbf{EC}_{c,j}\|}\right\rangle,$ (9)
where $j\in\{1,2,\dots,n_{dsc}\}$. To mitigate the risk of overfitting in the CLIP’s visual encoder, visual features with low relevance to either class text or GA descriptions should be filtered out, as they are unlikely to support meaningful visual-GA instance matching. Retaining these irrelevant features would introduce noise into the optimization directions during training, thereby increasing the risk of overfitting and accelerating knowledge forgetting. Hence, a condition $\chi(\mathbf{z}_{i})$ is defined to filter visual features as:
@@
We posit that embedding candidates in $\mathbf{EC}_{c}$ exhibiting a similarity score surpassing a threshold $\gamma$ ($\gamma>0$, relative to the anchor threshold $AT_{i}^{c}$) are more consistent with the visual features of the current sample. These candidates are filtered as follows:
@@
where $j\in\{1,2,\dots,n_{dsc}\}$, and sorted with descending order according to $\textbf{EC\_S}_{c}^{i,j}$. To reduce the potential influence of domain discrepancies arising from contextual descriptions, we further restrict the selection process by exclusively retaining attribute description sentences in $\mathbf{DC}_{c}$ that explicitly include the class name as a noun (i.e., <class name>+<GA>]).
@@ III-E General Attribute-Guided Progressive Visual-Textual Alignment
1) Instance Matching: To align the visual features with highly relevant embedded GA descriptions, we select the closest textual representation $\mathbf{h}_{i}$ as the paired text embedding:
@@
$\mathcal{L}_{\text{IM}}=\mathbb{E}_{i\sim P}\Big{[}-\log\frac{\mathbf{M}_{i,i}}{\mathbf{M}_{i,i}+\sum_{j\in P,j\neq i}\mathbf{M}_{i,j}}\Big{]},$ (14)
@@
$\mathcal{L}_{\text{TA}}^{i}=\mathbb{E}_{\mathbf{u}\sim\mathbf{FE}^{i}}\left\|\frac{\mathbf{w}^{\prime}_{t,k}}{\|\mathbf{w}^{\prime}_{t,k}\|}-\frac{\mathbf{u}}{\|\mathbf{u}\|}\right\|_{2}^{2}.$ (19)
Thus, text alignment loss $\mathcal{L}_{\text{TA}}$ of the current batch can be obtained by: $\mathcal{L}_{\text{TA}}=\sum_{i\sim P}\mathcal{L}_{\text{TA}}^{i}.$ Since, in the context of continual learning, data from previous tasks cannot be revisited during the current task, the weights $\mathbf{w^{\prime}}_{t}=\{\mathbf{w^{\prime}}_{t,1},\mathbf{w^{\prime}}_{t,2},...,\mathbf{w^{\prime}}_{t,|C_{t}|}\}$ are calibrated solely within the scope of the current task $t$. Consequently, only the shifting weights $\mathbf{s}_{t}=\{\mathbf{s}_{t,1},\mathbf{s}_{t,2},...,\mathbf{s}_{t,|C_{t}|}\}$ associated with the current task are learnable, whereas the shifting weights $\mathbf{s}_{0:t-1}=\{\mathbf{s}_{0},\mathbf{s}_{1},...,\mathbf{s}_{t-1}\}$ from previous tasks remain fixed.
@@
$\mathcal{L}_{\text{RIC}}=\mathbb{E}_{i\sim\{1,2,...,B\}}\Bigg{[}-\log\frac{\exp\Big{(}\left\langle\frac{\mathbf{z}_{i}}{\|\mathbf{z}_{i}\|},\frac{\mathbf{w^{\prime}}_{y_{i}}}{\|\mathbf{w^{\prime}}_{y_{i}}\|}\right\rangle/\tau\Big{)}}{\sum_{k=1}^{|C_{t}|}\exp\Big{(}\left\langle\frac{\mathbf{z}_{i}}{\|\mathbf{z}_{i}\|},\frac{\mathbf{w^{\prime}}_{t,k}}{\|\mathbf{w^{\prime}}_{t,k}\|}\right\rangle/\tau\Big{)}}\Bigg{]},$ (20)
@@ III-F Inference Stage
The GA descriptions and embeddings we introduce do not participate in the inference stage of the CLIP after each training phase, which effectively avoids additional storage overhead and eliminates any increase in inference time. Consequently, the model obtained through our method maintains identical parameter size and inference time during the inference stage as the original zero-shot CLIP. In inference stage after task $t$, we leverage the calibrated text embeddings of all seen classes: $\mathbf{W^{\prime}}_{t}=\bigcup_{j=1}^{t}\bigcup_{k=1}^{|C_{j}|}\mathbf{w^{\prime}}_{j,k}.$ Hence, the probability of predicting the test image $x$ as the class $y_{i}$ can be expressed as:
$p\left(y_{i}\mid x\right)=\frac{\exp\left(\left\langle\frac{\mathbf{z}}{\|\mathbf{z}\|},\frac{\mathbf{w^{\prime}}_{y_{i}}}{\|\mathbf{w^{\prime}}_{y_{i}}\|}\right\rangle/\tau\right)}{\sum_{\mathbf{w}^{\prime}_{j,k}\sim\mathbf{W^{\prime}}_{t}}\exp\left(\left\langle\frac{\mathbf{z}}{\|\mathbf{z}\|},\frac{\mathbf{w}^{\prime}_{j,k}}{\|\mathbf{w}^{\prime}_{j,k}\|}\right\rangle/\tau\right)}.$ (22)
@@ IV-A Setup
1) Datasets: The evaluation experiments for continual learning are conducted on CIFAR100 [66], ImageNet-Sub [24], CUB-200 [67], and ImageNet-R [68]. To further evaluate generalization under ‘unseen’ conditions, we additionally incorporate two out-of-distribution (OOD) datasets: FGVC-Aircraft [69] and SKIN-40 [70].
@@
2) Metrics: To evaluate the continual learning performance of classification models, we employ two primary metrics: ‘Last’ and ‘Avg’. ‘Last’ denotes the average accuracy across all classes after the model has completed training on the final task. ‘Avg’ represents the mean incremental accuracy calculated over all tasks the model has learned thus far. In addition, the control set accuracy ‘C.’ [2] is employed to evaluate the retention of CLIP’s zero-shot generalization knowledge after continual learning, assessed on ImageNet-1K [24] (1000 classes in total).
TABLE I: Comparison of different methods on the CIFAR100, ImageNet-Sub, and CUB-200 datasets under various settings with CLIP ViT-L/14. ‘UB’ denotes the upper bound achieved through joint training.
Method Mem. of Old Tasks CIFAR100 ImageNet-Sub CUB-200
$T=5$ $T=10$ $T=20$ $T=5$ $T=10$ $T=20$ $T=5$ $T=10$ $T=20$
Seq FT 50.6 66.9 58.8 73.8 40.8 64.6
LwF [12] 56.1 74.8 56.4 75.0 52.1 71.5
EWC [32] Fisher Approx. 64.4 78.4 70.3 80.2 64.4 78.6
AdaBOP [41] Principal Statistics 82.1 86.7 81.0 83.9 71.2 80.0
OSN [42] Principal Statistics 83.6 88.7 82.6 84.5 73.2 80.9
ContinualCLIP [64] 72.8 72.8 72.8 74.6 74.6 74.6 60.9 60.9 60.9
L2P [45] 70.2 79.6 71.1 78.4 69.9 77.8
DualPrompt [45] Task-aware Key-prompt 72.0 81.8 71.7 79.8 64.5 75.9
SLCA [1] Class-wise Statistics 81.8 87.3 80.2 86.6 79.2 87.8 82.3 86.0 80.3 84.4 80.2 85.4 75.5 78.9 73.9 80.0 71.1 79.4
ENGINE [71] Class-wise Statistics 81.0 85.9 79.9 86.0 77.3 84.0 76.3 82.0 74.0 81.2 68.9 76.7
ENGINE(w/ GDA) [71] Class-wise Statistics 83.1 87.9 81.5 87.7 79.9 86.9 78.8 82.5 78.9 83.4 73.7 80.3
UB 90.0 90.0 90.0 87.6 87.6 87.6 84.5 84.5 84.5
@@
3) Competitors: We compare our method against baseline and state-of-the-art methods for continual learning. These include ContinualCLIP [64] under zero-shot conditions; non-continual adapting methods such as CoOp [5] and CoOpOp [6]; and conventional continual learning method such as sequential fine-tuning (Seq FT), LwF [12], EWC [32], and recent task-orthogonal methods: OSN [42] and AdaBOP [41]. Additionally, we evaluate VLM-specific continual learning methods, including AttriCLIP [16], TaskRes-CL [7], SPU [2], MoE-Adapter [18], RAPF [72], and ENGINE [71](‘w/ GDA’ represents using Gaussian Discriminant Analysis to correct the logits). For a broader comparison, we also incorporate techniques tailored to address continual learning in visual-only pre-trained models, such as L2P [45], DualPrompt [46], and SLCA [1].
@@
4) Implementation Details: All experiments are conducted on two NVIDIA GeForce RTX 3090 GPUs. We adopt the pre-trained CLIP [3] ViT-L/14 and ViT-B/16 as the backbones. To ensure a fair comparison, all methods are built upon the same pre-trained CLIP backbones, and no data replay strategies are used during the continual learning process.
@@
The model is trained for 10 epochs on the datasets for each incremental task. Stochastic Gradient Descent (SGD) [73] is employed as the optimizer, using a cosine learning rate decay schedule with a batch size of 32. The learning rate for the MLPs’ weights in the CLIP’s visual encoder is set to $1\times 10^{-5}$, while the learning rate for the text shifting weights is set to 0.1. The coefficient $\alpha$ for shifting weight addition is set to 0.1. Loss balancing factors are set as $\lambda_{\text{IM}}=2.0$, $\lambda_{\text{TA}}=2.0$, and $\lambda_{\text{RIC}}=1.0$ by default. For the fine-grained CUB-200 and FGVC-Aircraft, $\lambda_{\text{IM}}$ is adjusted to 15.0, and 5.0. The threshold parameters $\delta_{d}$ and $\gamma$ are first estimated based on downstream tasks, and then set based on empirical observations. The specific estimation procedure and detailed analysis are presented in IV-C-2).
Additionally, we generate $n_{dsc}=30$ attribute descriptions for each class using the language assistant GPT-4 [54].